# Nonuniform Sobolev Spaces

### 05 Jan 2024 v1

@@ Abstract.
We study nonuniform Sobolev spaces, i.e., spaces of functions whose partial derivatives lie in possibly different Lebesgue spaces. Although standard proofs do not apply, we show that nonuniform Sobolev spaces share similar properties as the classical ones. These spaces arise naturally in the study of certain PDEs. For instance, we illustrate that nonuniform fractional Sobolev spaces are useful in the study of convergence of Schrödinger operators. In this work we extend recent advances on the convergence of Schrödinger operators to nonuniform fractional Sobolev spaces.
@@
Sobolev spaces, Nonuniform Sobolev spaces, fractinoal Sobolev spaces, embedding theorems, Schrödinger operators.
@@ 1. Introduction
The theory of Sobolev spaces plays an import role in the study of partial differential equations and many other fields. Given a positive integer $m$, a positive number $p\in[1,\infty]$ and an open set $\Omega\subset\mathbb{R}^{N}$, we denote by
@@
the classical Sobolev space, where $\alpha=(\alpha_{1},\ldots,\alpha_{N})$ is a multi-index, $|\alpha|=\sum_{i=1}^{N}\alpha_{i}$, and $D^{\alpha}$ denotes the weak partial derivatives, that is, partial derivatives in the sense of distributions.
@@
For the nonuniform case, that is, for $f\in C^{1}\cap W_{k}^{\vec{p}}$ with $\vec{p}=(p_{0},p_{1})$, both (1.2) and (1.3) are still true. However, their proofs are quite different from the classical uniform case. In the classical case, one first obtains the density of compactly supported differentiable functions, then derives (1.2) from (1.3) for such functions, and finally extends (1.2) to all functions in $W^{1,p_{1}}$ by density. For the nonuniform case, compactly supported differentiable functions are still dense in $W_{1}^{\vec{p}}$. However, the embedding inequality is required in the proof. So we have to prove (1.2) directly for functions which is not compactly supported. In this case, (1.2) is not a straightforward consequence of (1.3): we have to show first that $f\in L^{Np_{1}/(N-p_{1})}$.
Another application of nonuniform Sobolev spaces is to study the convergence of Schrödinger operators defined by
@@
Since then, many works have been done on this topic. Dahlberg and Kenig [14] provided counterexamples indicating that the range $s\geq 1/4$ is sharp for $N=1$. And Sjölin [31] extended this result to the case $a>1$.
@@
Related works also include the non-tangential convergence by Shiraki [30], Yuan, Zhao and Zheng [36], Li, Wang and Yan [24, 25], the convergence along curves by Cao and Miao [8] and Zheng [37], and the convergence rate by Cao, Fan and Wang [7]; see also the works by Cowling [13], Walther [35], and Rogers and Villarroya [27] for the case $a<1$.
@@
is the Gagliardo seminorm of $f$. Fractional Sobolev spaces were introduced by Aronszajn [2], Gagliardo [20] and Slobodeckiĭ [32]. We refer to [1, 5, 17] for an overview of fractional Sobolev spaces and [6, 22] for the limit of norms of fractional Sobolev spaces.
@@
The paper is organized as follows. In Section 2, we show that compactly supported infinitely differentiable functions are dense in nonuniform Sobolev spaces. We also obtain the Sobolev inequality for functions in nonuniform Sobolev spaces. In Section 3, we focus on nonuniform fractional Sobolev spaces and present an embedding theorem for such spaces. And in Section 4, we give an application of nonuniform Sobolev spaces. We prove the almost everywhere convergence of Schrödinger operators for a large class of functions, extending known results.
@@ 2. Embedding Inequalities for Nonuniform Sobolev Spaces
Take some $\psi\in C_{c}^{\infty}$ such that $\psi(x)=1$ whenever $|x|<1$. Set $g_{n}(x)=\psi(x/n)g(x)$.
It remains to show that the sequence $\{g_{n}:\,n\geq 1\}$ converges to $g$ in $W_{k}^{\vec{p}}$.
@@
    If $kp_{k}>N$, then
    $W_{k}^{\vec{p}}(\mathbb{R}^{N})\hookrightarrow L^{\infty}(\mathbb{R}^{N}).$
@@
(iii)   Finally, we consider the case $kp_{k}>N$. We prove it by induction on $k$.
For $k=1$, the arguments in [16, pages 80-82] work well for the nonuniform case with minor changes. Specifically, applying the $L^{p_{1}}$ norm for derivatives and the $L^{p_{0}}$ norm for the function itself, the same arguments yield that $W_{1}^{\vec{p}}\hookrightarrow L^{\infty}$.
Now we assume that the conclusion is true for the cases $1$, $\ldots$, $k-1$. Consider the case $k$. Fix some $f\in W_{k}^{\vec{p}}$. There are two cases.
@@
Consequently, $f\in W_{1}^{(p_{0},q)}$. Applying the inductive assumption for the case $k=1$, we get
@@
In this case, we see from (i) that $f\in W_{1}^{(p_{0},q_{1})}$ with $q_{1}=Np_{k}/(N-(k-1)p_{k})$. Since $kp_{k}>N$, we have $q_{1}>N$. Applying the inductive assumption for the case $k=1$ again, we get
@@
By induction, the embedding $\|f\|_{L^{\infty}}\lesssim\|f\|_{W_{k}^{\vec{p}}}$ is valid for all $k\geq 1$. ∎
Theorem 2.7 has some interesting consequences. For example, the following result shows that the Sobolev space $W_{k}^{\vec{q}}$ with $1\leq q_{k}<N/k$ and $1/q_{i}=1/q_{k}-(k-i)/N$ is the largest one among all Sobolev spaces $W_{k}^{\vec{p}}$ with $p_{k}=q_{k}$.
@@ 3.1. Embedding Theorem for Nonuniform Fractional Sobolev Spaces
    If $sp_{1}>N$, then $W_{s}^{\vec{p}}(\mathbb{R}^{N})\hookrightarrow L^{\infty}(\mathbb{R}^{N})$.
@@
Recall that for a tempered distribution $\Lambda$ and a function $h\in C^{\infty}$, for which the function itself and all of its derivatives have at most polynomial growth at infinity, the product $h\Lambda$ is the tempered distribution [21, Definition 2.3.15] defined by
@@
Since $g(0,x)=f(x)$ a.e., it suffices to show that $\|g(0,\cdot)\|_{L^{p_{1}}}\lesssim\|f\|_{W_{s}^{\vec{p}}}$. We prove it in the following steps.
@@
(S3)  Now we deal with the last two terms.
Observe that both $\nabla\Phi$ and $\Delta\Phi$ vanish in a neighborhood of 00. Hence both $(\nabla\Phi\cdot\nabla E)$ and $(\Delta\Phi)E$ are compactly supported bounded functions. Hence
@@
Suppose that $0<s\leq 1$, $1\leq p_{0},p_{1}<\infty$ and $p_{1}\alpha<N<p_{0}\alpha$. Let $f(x)=(1+|x|^{2})^{-\alpha/2}$, $x\in\mathbb{R}^{N}$ and let $\vec{p}=(p_{0},p_{1})$. We have $f\in W_{s}^{\vec{p}}(\mathbb{R}^{N})$ if and only if $p_{1}(\alpha+s)>N$.
Consequently, for any $0<s\leq 1$ and $p_{0}>p_{1}$, $W_{s}^{\vec{p}}\not\subset W_{s}^{p_{1}}$.
@@
(i)  We show that $f\in W_{s}^{\vec{p}}(\mathbb{R}^{N})$ if and only if $p_{1}(\alpha+s)>N$, for which we only need to consider the case $10<s<$ 1 since the other case $s=1$ is obvious.
First, we assume that $p_{1}(\alpha+s)>N$. Let us estimate the integral
@@
$\lesssim|a|^{p_{1}-1}\bigg{|}\int_{0}^{a}\int_{|x|>2|a|}\frac{|x_{j}+t|^{p_{1}}}{(1+|\hat{x}_{j}|^{2}+|x_{j}+t|^{2})^{(\alpha+2)p_{1}/2}}\mathrm{d}x\,\mathrm{d}t\bigg{|}$
$\leq|a|^{p_{1}-1}\bigg{|}\int_{0}^{a}\int_{|x|>|a|}\frac{|x_{j}|^{p_{1}}}{(1+|\hat{x}_{j}|^{2}+|x_{j}|^{2})^{(\alpha+2)p_{1}/2}}\mathrm{d}x\,\mathrm{d}t\bigg{|}$
$=|a|^{p_{1}}\bigg{(}\int_{|x_{j}|>|a|}\int_{\mathbb{R}^{N-1}}\frac{|x_{j}|^{p_{1}}}{(1+|\hat{x}_{j}|^{2}+|x_{j}|^{2})^{(\alpha+2)p_{1}/2}}\mathrm{d}\hat{x}_{j}\,\mathrm{d}x_{j}$
$\qquad+\int_{|x_{j}|<|a|}\int_{|\hat{x}_{j}|>(a^{2}-x_{j}^{2})^{1/2}}\frac{|x_{j}|^{p_{1}}}{(1+|\hat{x}_{j}|^{2}+|x_{j}|^{2})^{(\alpha+2)p_{1}/2}}\mathrm{d}\hat{x}_{j}\,\mathrm{d}x_{j}\bigg{)}$
$\approx|a|^{p_{1}}\bigg{(}\int_{|x_{j}|>|a|}\frac{|x_{j}|^{p_{1}}\mathrm{d}x_{j}}{(1+|x_{j}|)^{(\alpha+2)p_{1}-(N-1)}}+\int_{|x_{j}|<|a|}\frac{|x_{j}|^{p_{1}}\mathrm{d}x_{j}}{(1+|a|)^{(\alpha+2)p_{1}-(N-1)}}\bigg{)}$
$\approx|a|^{N-p_{1}\alpha},$ (3.15)
where we use the fact that $(\alpha+1)p_{1}>(\alpha+s)p_{1}>N$. On the other hand,
@@
$\lesssim\int_{|x|\leq 2|a|}\Big{(}\frac{1}{(1+|x|^{2})^{p_{1}\alpha/2}}+\frac{1}{(1+|x+ae_{j}|^{2})^{p_{1}\alpha/2}}\Big{)}\mathrm{d}x$
$\leq\int_{|x|\leq 2|a|}\frac{1}{(1+|x|^{2})^{p_{1}\alpha/2}}\mathrm{d}x+\int_{|x+ae_{j}|\leq 3|a|}\frac{1}{(1+|x+ae_{j}|^{2})^{p_{1}\alpha/2}}\mathrm{d}x$
$\lesssim|a|^{N-p_{1}\alpha}.$ (3.16)
@@
$\int_{\mathbb{R}^{N}}|f(x)-f(x+ae_{j})|^{p_{1}}\mathrm{d}x\lesssim|a|^{N-p_{1}\alpha},\quad|a|>1.$
For $|a|\leq 1$, since $|\partial_{x_{j}}f(x)|\lesssim 1/(1+|x|^{2})^{(\alpha+1)/2}$ and $p_{1}(\alpha+1)>p_{1}(\alpha+s)>N$, we have $\partial_{x_{j}}f\in L^{p_{1}}$. It follows from (3.14) that
@@
$\lesssim\int_{|a|\leq 1}|a|^{p_{1}-1-p_{1}s}\mathrm{d}a+\int_{|a|>1}\frac{\mathrm{d}a}{|a|^{p_{1}(\alpha+s)-N+1}}$
@@
Next we show that $f\not\in W_{s}^{\vec{p}}$ whenever $p_{1}(\alpha+s)\leq N$.
@@
$\gtrsim\int_{\begin{subarray}{c}a^{2}<|x|^{2}<5a^{2}/4\\ x_{j}>0\end{subarray}}\frac{\mathrm{d}x}{a^{p_{1}\alpha}}$
$\approx a^{N-p_{1}\alpha}.$

### 23 Jan 2024 v3

@@ Abstract.
We study nonuniform Sobolev spaces, i.e., spaces of functions whose partial derivatives lie in possibly different Lebesgue spaces. Although standard proofs do not apply, we show that nonuniform Sobolev spaces share similar properties as the classical ones. These spaces arise naturally in the study of certain PDEs. For instance, we illustrate that nonuniform fractional Sobolev spaces are useful in the study of local estimates for solutions of heat equations and the convergence of Schrödinger operators. In this work we extend recent advances on local energy estimates for solutions of heat equations and the convergence of Schrödinger operators to nonuniform fractional Sobolev spaces.
@@
Sobolev spaces, Nonuniform Sobolev spaces, fractinoal Sobolev spaces, embedding theorems, heat equations, Schrödinger operators.
@@ 1. Introduction
Given a positive integer $m$, a positive number $p\in[1,\infty]$ and an open set $\Omega\subset\mathbb{R}^{N}$, we denote by
@@
the classical Sobolev space, where $\alpha=(\alpha_{1},\ldots,\alpha_{N})$ is a multi-index, $|\alpha|=\sum_{i=1}^{N}\alpha_{i}$, and $D^{\alpha}$ denotes the weak partial derivatives, that is, partial derivatives in the sense of distributions. The theory of Sobolev spaces plays an import role in the study of partial differential equations and many other fields. We refer to text books by Adams and Fournier [1], Demengel and Demengel [16], Evans [20], Grafakos [25, 26], Leoni [29], Pişkin and Okutmuştur [34] for an overview of Sobolev spaces and applications in PDEs and in harmonic analysis.
@@
For the nonuniform case, that is, for $f\in C^{1}\cap W_{k}^{\vec{p}}$ with $\vec{p}=(p_{0},p_{1})$, both (1.2) and (1.3) are still true. However, their proofs are quite different from the uniform case. In the classical uniform case, one first obtains the density of compactly supported differentiable functions, then derives (1.2) from (1.3) for such functions, and finally extends (1.2) to all functions in $W^{1,p_{1}}$ by density. For the nonuniform case, compactly supported differentiable functions are still dense in $W_{1}^{\vec{p}}$. However, the embedding inequality is required in the proof. So we have to prove (1.2) directly for functions which is not compactly supported. In this case, (1.2) is not a straightforward consequence of (1.3): we have to show first that $f\in L^{Np_{1}/(N-p_{1})}$.
In some applications, only the derivatives are concerned. For example, Fefferman, Israel and Luli [21, 22] studied Sobolev extension operators for homogeneous Sobolev spaces. In this case, it is natural to consider nonuniform Sobolev spaces.
Another application we discuss concerns local estimates for solutions of heat equations. We obtain local energy estimates for initial data in nonuniform Sobolev spaces; this extends the local estimates by Fefferman, McCormick, Robinson and Rodrigo [23].
Moreover, nonuniform Sobolev spaces are also useful in the study of the convergence of Schrödinger operators defined by
@@
Since then, many works have appeared on this topic. Dahlberg and Kenig [14] provided counterexamples indicating that the range $s\geq 1/4$ is sharp for $N=1$. And Sjölin [40] extended this result to the case $a>1$.
@@
Related works also include the non-tangential convergence by Shiraki [39], Yuan, Zhao and Zheng [45], Li, Wang and Yan [32, 33], the convergence along curves by Cao and Miao [8] and Zheng [46], the Hausdorff dimension of the divergence set by Li, Li and Xiao [31], and the convergence rate by Cao, Fan and Wang [7]; see also the works by Cowling [13], Walther [44], and Rogers and Villarroya [36] for the case $a<1$.
@@
is the Gagliardo seminorm of $f$. Fractional Sobolev spaces were introduced by Aronszajn [2], Gagliardo [24] and Slobodeckiĭ [41]. We refer to the text book by Leoni [30] and papers by Brezis and Mironescu [5] and Di Nezza [17] for an overview of fractional Sobolev spaces. See also recent papers by Brezis, Van Schaftingen and Yung [6] and Gu and Yung [27] for the limit of norms of fractional Sobolev spaces.
@@
The paper is organized as follows. In Section 2, we show that compactly supported infinitely differentiable functions are dense in nonuniform Sobolev spaces. We also obtain the Sobolev inequality for functions in nonuniform Sobolev spaces. In Section 3, we focus on nonuniform fractional Sobolev spaces and present an embedding theorem for such spaces. And in Section 4, we give two applications of nonuniform Sobolev spaces. We give local energy estimates for solutions of the heat equation with initial data in nonuniform Sobolev spaces. And we prove the almost everywhere convergence of Schrödinger operators for a large class of functions, extending known results.
@@ 2. Embedding Inequalities for Nonuniform Sobolev Spaces
Take some $\psi\in C_{c}^{\infty}$ such that $\psi(x)=1$ whenever $|x|<1$. Set $g_{n}(x)=\psi(x/n)g(x)$. It remains to show that the sequence $\{g_{n}:\,n\geq 1\}$ converges to $g$ in $W_{k}^{\vec{p}}$.
@@
Recall that for an integer $n\geq 0$, the space $C_{b}^{n}(\mathbb{R}^{N})$ consists of all functions $f$ such that $f$ is $n$ times continuously differentiable and for any multi-index $\alpha$ with $|\alpha|\leq n$, $D^{\alpha}f\in L^{\infty}$.
For an integer $n\geq 0$ and a positive number $\nu\in(0,1)$, the space $C^{n,\nu}(\mathbb{R}^{N})$ consists of all functions $f$ in $C_{b}^{n}$ such that for any multi-index $\alpha$ with $|\alpha|=n$,
$|D^{\alpha}f(x)-D^{\alpha}f(y)|\leq C_{n,\nu}|x-y|^{\nu},\quad\forall x,y\in\mathbb{R}^{N}.$
@@
    If $kp_{k}>N$, then $W_{k}^{\vec{p}}(\mathbb{R}^{N})\hookrightarrow L^{\infty}(\mathbb{R}^{N})$. More precisely, if $kp_{k}>N$ and $N/p_{k}\not\in\mathbb{Z}$, then there is some integer $k_{0}$ such that $(k_{0}-1)p_{k}<N<k_{0}p_{k}$ and
    $W_{k}^{\vec{p}}(\mathbb{R}^{N})\hookrightarrow C_{b}^{k-k_{0},k_{0}-N/p_{k}}(\mathbb{R}^{N}).$
    If $N/p_{k}\in\mathbb{Z}$ and $k\geq k_{0}:=N/p_{k}+1$, then for any $10<\lambda<$ 1,
    $W_{k}^{\vec{p}}(\mathbb{R}^{N})\hookrightarrow C_{b}^{k-k_{0},\lambda}(\mathbb{R}^{N}).$
@@
(iii)   Finally, we consider the case $kp_{k}>N$.
First, we show that $W_{k}^{\vec{p}}\subset C_{b}^{0}$. We prove it by induction on $k$. For $k=1$, the arguments in [16, pages 80-82] work well for the nonuniform case with minor changes. Specifically, applying the $L^{p_{1}}$ norm for derivatives and the $L^{p_{0}}$ norm for the function itself, the same arguments yield that $W_{1}^{\vec{p}}\hookrightarrow C_{b}^{0}$. That is, functions in $W_{1}^{\vec{p}}$ are continuous and bounded.
Now we assume that the conclusion is true for the cases $1$, $\ldots$, $k-1$. Consider the case $k$. Fix some $f\in W_{k}^{\vec{p}}$. There are two cases:
@@
Consequently, $f\in W_{1}^{(p_{0},q)}$. Applying the inductive assumption for the case $k=1$, we get that $f$ is continuous and
@@
In this case, we see from (i) that $f\in W_{1}^{(p_{0},q_{1})}$ with $q_{1}=Np_{k}/(N-(k-1)p_{k})$. Since $kp_{k}>N$, we have $q_{1}>N$. Applying the inductive assumption for the case $k=1$ again, we get that $f$ is continuous and
@@
By induction, the embedding $W_{k}^{\vec{p}}\hookrightarrow C_{b}^{0}$ is valid for all $k\geq 1$.
Next we prove the Hölder continuity.
First, we assume that $N/p_{k}$ is not an integer. When $k=1$, we have $p_{1}>N$. Take some $f\in W_{1}^{\vec{p}}$. Let $\varphi$ and $g_{\lambda}$ be defined as in Lemma 2.1. We have
$g_{\lambda}(x+y)-g_{\lambda}(x)=\int_{0}^{1}\nabla g_{\lambda}(x+ty)\cdot y\mathrm{d}t=\int_{0}^{1}\int_{\mathbb{R}^{N}}\nabla f(x+ty-z)\cdot y\varphi(z)\mathrm{d}z\mathrm{d}t,\quad\forall x,y\in\mathbb{R}^{N}.$
It follows from Minkowski’s inequality that
$\|g_{\lambda}(\cdot+y)-g_{\lambda}\|_{L^{p_{1}}}\leq|y|\cdot\|\nabla f\|_{L^{p_{1}}}.$
Since $f$ is continuous and $\lim_{\lambda\rightarrow 0}g_{\lambda}(x)=f(x)$ for all $x\in\mathbb{R}^{N}$, letting $\lambda\rightarrow 0$ in the above inequality, we see from Fatou’s lemma that
$\|f(\cdot+y)-f\|_{L^{p_{1}}}\leq|y|\cdot\|\nabla f\|_{L^{p_{1}}}.$
On the other hand, it is easy to see that
$\|\nabla(f(\cdot+y)-f)\|_{L^{p_{1}}}\leq 2\|\nabla f\|_{L^{p_{1}}}.$
Applying the fact for classical Sobolev spaces that if $u,|\nabla u|\in L^{p_{1}}$ and $p_{1}>N$, then
$\|u\|_{L^{\infty}}\lesssim\|u\|_{L^{p_{1}}}^{1-N/p_{1}}\|\nabla u\|_{L^{p_{1}}}^{N/p_{1}},$ (2.14)
we get
$\|f(\cdot+y)-f\|_{L^{\infty}}\leq|y|^{1-N/p_{1}}\cdot\|\nabla f\|_{L^{p_{1}}}.$
Hence $f\in C_{b}^{0,1-N/p_{1}}$.
When $k\geq 2$, there is some positive integer $k_{0}\leq k$ such that $(k_{0}-1)p_{k}<N<k_{0}p_{k}$. It follows that for any multi-index $\alpha$ with $|\alpha|=k-k_{0}$, we have $D^{\alpha}f\in W_{k_{0}}^{(p_{k-k_{0}},\ldots,p_{k})}$. Now we see from (i) that for any multi-index $\beta$ with $|\beta|=1$,
$D^{\alpha+\beta}f\in L^{Np_{k}/(N-(k_{0}-1)p_{k})}.$
Since $Np_{k}/(N-(k_{0}-1)p_{k})>N$, applying the conclusion for $k=1$, we get $D^{\alpha}f\in C_{b}^{0,1-N/(Np_{k}/(N-(k_{0}-1)p_{k}))}=C_{b}^{0,k_{0}-N/p_{k}}$.
Moreover, by the embedding inequality we have proved, $D^{\beta}f\in L^{q}$ whenever $1\leq|\beta|\leq k-k_{0}$ and $q$ is large enough. In particular, $D^{\beta}f\in L^{q}$ for some $q>N$. Applying the conclusion for $k=1$ again, we get $D^{\alpha}f\in C_{b}^{0}$ when $|\alpha|\leq k-k_{0}-1$. Hence $f\in C_{b}^{k-k_{0},k_{0}-N/p_{k}}$.
It remains to consider the case $N/p_{k}\in\mathbb{Z}$. When $(k_{0}-1)=N<k_{0}p_{k}$, for any multi-index $\alpha$ with $|\alpha|=k-k_{0}$, we have $D^{\alpha}f\in W_{k_{0}}^{(p_{k-k_{0}},\ldots,p_{k})}$. Now we see from (ii) that for $q$ large enough and $|\beta|=1$,
$D^{\alpha+\beta}f\in L^{q}.$
The arguments for the case $k=1$ show that for $q$ large enough, $D^{\alpha}f\in C_{b}^{0,1-N/q}$. Consequently, $D^{\alpha}f\in C_{b}^{0,\lambda}$ whenever $|\alpha|=k-k_{0}$ and $10<\lambda<$ 1.
On the other hand, the same arguments as those for the case $N/p_{k}\not\in\mathbb{Z}$ show that $D^{\alpha}f\in C_{b}^{0}$ when $|\alpha|\leq k-k_{0}-1$. Hence $f\in C_{b}^{k-k_{0},\lambda}$ for any $10<\lambda<$ 1. ∎
Theorem 2.7 has some interesting consequences. In fact, the following corollary can be proved with similar arguments as in the proof of Theorem 2.7 (iii), for which we leave the details to interested readers.
Corollary 2.8.
Let $k$ be a positive integer and $\vec{p}=(p_{0},\ldots,p_{k})\in[1,\infty)^{k+1}$. If $p_{k}>N$, then
$W_{k}^{\vec{p}}\hookrightarrow C_{b}^{k-1,1-N/p_{k}}.$
On the other hand, the following result shows that the Sobolev space $W_{k}^{\vec{q}}$ with $1\leq q_{k}<N/k$ and $1/q_{i}=1/q_{k}-(k-i)/N$ is the largest one among all Sobolev spaces $W_{k}^{\vec{p}}$ with $p_{k}=q_{k}$.
@@ 3.1. Embedding Theorem for Nonuniform Fractional Sobolev Spaces
    If $sp_{1}>N$, then $W_{s}^{\vec{p}}(\mathbb{R}^{N})\hookrightarrow C_{b}^{0,s-N/p_{1}}(\mathbb{R}^{N})$.
@@
Recall that for a tempered distribution $\Lambda$ and a function $h\in C^{\infty}$, for which the function itself and all of its derivatives have at most polynomial growth at infinity, the product $h\Lambda$ is the tempered distribution [25, Definition 2.3.15] defined by
@@
Since $g(0,x)=f(x)$ a.e., it suffices to show that $\|g(0,\cdot)\|_{L^{p_{1}}}\lesssim\|f\|_{W_{s}^{\vec{p}}}$. We prove it in the following steps:
@@
(S3)  We deal with the last two terms. Observe that both $\nabla\Phi$ and $\Delta\Phi$ vanish in a neighborhood of 00. Hence both $(\nabla\Phi\cdot\nabla E)$ and $(\Delta\Phi)E$ are compactly supported bounded functions. Hence
@@
Suppose that $0<s\leq 1$, $1\leq p_{0},p_{1}<\infty$ and $p_{1}\delta<N<p_{0}\delta$. Let $f(x)=(1+|x|^{2})^{-\delta/2}$, $x\in\mathbb{R}^{N}$ and let $\vec{p}=(p_{0},p_{1})$. We have $f\in W_{s}^{\vec{p}}(\mathbb{R}^{N})$ if and only if $p_{1}(\delta+s)>N$. Consequently, for any $0<s\leq 1$ and $p_{0}>p_{1}$, $W_{s}^{\vec{p}}\not\subset W_{s}^{p_{1}}$.
@@
(i)  We show that $f\in W_{s}^{\vec{p}}(\mathbb{R}^{N})$ if and only if $p_{1}(\delta+s)>N$, for which we only need to consider the case $10<s<$ 1 since the other case $s=1$ is obvious.
First, we assume that $p_{1}(\delta+s)>N$. Let us estimate the integral
@@
$\lesssim|a|^{p_{1}-1}\bigg{|}\int_{0}^{a}\int_{|x|>2|a|}\frac{|x_{j}+t|^{p_{1}}}{(1+|\hat{x}_{j}|^{2}+|x_{j}+t|^{2})^{(\delta+2)p_{1}/2}}\mathrm{d}x\,\mathrm{d}t\bigg{|}$
$\leq|a|^{p_{1}-1}\bigg{|}\int_{0}^{a}\int_{|x|>|a|}\frac{|x_{j}|^{p_{1}}}{(1+|\hat{x}_{j}|^{2}+|x_{j}|^{2})^{(\delta+2)p_{1}/2}}\mathrm{d}x\,\mathrm{d}t\bigg{|}$
$=|a|^{p_{1}}\bigg{(}\int_{|x_{j}|>|a|}\int_{\mathbb{R}^{N-1}}\frac{|x_{j}|^{p_{1}}}{(1+|\hat{x}_{j}|^{2}+|x_{j}|^{2})^{(\delta+2)p_{1}/2}}\mathrm{d}\hat{x}_{j}\,\mathrm{d}x_{j}$
$\qquad+\int_{|x_{j}|<|a|}\int_{|\hat{x}_{j}|>(a^{2}-x_{j}^{2})^{1/2}}\frac{|x_{j}|^{p_{1}}}{(1+|\hat{x}_{j}|^{2}+|x_{j}|^{2})^{(\delta+2)p_{1}/2}}\mathrm{d}\hat{x}_{j}\,\mathrm{d}x_{j}\bigg{)}$
$\approx|a|^{p_{1}}\bigg{(}\int_{|x_{j}|>|a|}\frac{|x_{j}|^{p_{1}}\mathrm{d}x_{j}}{(1+|x_{j}|)^{(\delta+2)p_{1}-(N-1)}}+\int_{|x_{j}|<|a|}\frac{|x_{j}|^{p_{1}}\mathrm{d}x_{j}}{(1+|a|)^{(\delta+2)p_{1}-(N-1)}}\bigg{)}$
$\approx|a|^{N-p_{1}\delta},$ (3.15)
where we use the fact that $(\delta+1)p_{1}>(\delta+s)p_{1}>N$. On the other hand,
@@
$\lesssim\int_{|x|\leq 2|a|}\Big{(}\frac{1}{(1+|x|^{2})^{p_{1}\delta/2}}+\frac{1}{(1+|x+ae_{j}|^{2})^{p_{1}\delta/2}}\Big{)}\mathrm{d}x$
$\leq\int_{|x|\leq 2|a|}\frac{1}{(1+|x|^{2})^{p_{1}\delta/2}}\mathrm{d}x+\int_{|x+ae_{j}|\leq 3|a|}\frac{1}{(1+|x+ae_{j}|^{2})^{p_{1}\delta/2}}\mathrm{d}x$
$\lesssim|a|^{N-p_{1}\delta}.$ (3.16)
@@
$\int_{\mathbb{R}^{N}}|f(x)-f(x+ae_{j})|^{p_{1}}\mathrm{d}x\lesssim|a|^{N-p_{1}\delta},\quad|a|>1.$
For $|a|\leq 1$, since $|\partial_{x_{j}}f(x)|\lesssim 1/(1+|x|^{2})^{(\delta+1)/2}$ and $p_{1}(\delta+1)>p_{1}(\delta+s)>N$, we have $\partial_{x_{j}}f\in L^{p_{1}}$. It follows from (3.14) that
@@
$\lesssim\int_{|a|\leq 1}|a|^{p_{1}-1-p_{1}s}\mathrm{d}a+\int_{|a|>1}\frac{\mathrm{d}a}{|a|^{p_{1}(\delta+s)-N+1}}$
@@
Next we show that $f\not\in W_{s}^{\vec{p}}$ whenever $p_{1}(\delta+s)\leq N$.
@@
$\gtrsim\int_{\begin{subarray}{c}a^{2}<|x|^{2}<5a^{2}/4\\ x_{j}>0\end{subarray}}\frac{\mathrm{d}x}{a^{p_{1}\delta}}$
$\approx a^{N-p_{1}\delta}.$